# Graph Learning

### v1

@@ 2.3.2 Mini-batch Training
Despite their effectiveness, typical graph learning methods face challenges when applied to dynamic or evolving graphs. Static models assume that the graph structure and node features do not change over time, which is often unrealistic in real-world scenarios where graph structures continuously evolve [barros2021survey_dynamicgraphsurvey, gravina2024deep_dynamicgraphSurveyBenchmark, febrinanto2023graph_GLL, zheng2025survey].
@@ 3.2 Spatiotemporal Graph Learning
$A_{u,v}=\frac{\exp\left(\frac{\log\widetilde{A}_{u,v}+\epsilon_{v}}{\tau}\right)}{\sum_{w\in\mathcal{N}_{u}}\exp\left(\frac{\log\widetilde{A}_{u,v}+\epsilon_{w}}{\tau}\right)},$ (3.2)
@@ 3.2.3 Attention-based Methods
Multimodal fusion aims to integrate information from diverse data sources, such as visual, textual, and acoustic data. However, multimodal data often exhibits intricate relational dependencies, including both inter-modal and intra-modal correlations. Traditional multimodal learning methods that rely on basic modality alignment and simple fusion techniques often fail to effectively capture and integrate the rich, complementary information across different modalities. To address this challenge, multimodal graph learning has emerged as a powerful approach [ektefaie2023multimodal, peng2024learning, xie2024multimodal, he2025unigraph2]. It leverages graph structures to model and learn the intricate inter-modal and intra-modal correlations, enabling a more nuanced integration of multimodal data.
@@ 4.1.1 Separate Unimodal Learning
Modality-specific Representation. Separate unimodal learning exhibits the advantage of capturing modality-specific representation. Specifically, the modality-specific representation $\mathbf{Z}_{i}$ for the modality $i$ can be effectively extracted using various independent graph learning models. For example, MGAT [tao2020mgat] applies embedding propagation layers to perform message passing on each single-modal user item graph, capturing user preferences on each modality. guo2024lgmrec propose local and global graph learning-guided multimodal recommender (LGMRec) for learning on multimodal user-item graphs. It introduces a local graph embedding module to perform message propagation on textual and visual user-item graphs separately, capturing intra-modal local user interests and generating modality-specific representations. MKGformer [chen2022hybrid] is a hybrid Transformer model designed for the multimodal knowledge graph completion task. In MKGformer, two types of knowledge graph are constructed: visual knowledge graphs and textual knowledge graphs. A visual Transformer and a textual Transformer, both equipped with self-attention modules, are then applied to these knowledge graphs to learn modality-specific representations.
@@
  • Contrastive learning-based fusion: Some works present multimodal graph contrastive learning frameworks. These frameworks apply a contrastive objective strategy to capture inter-modal and intra-modal similarities and differences, enabling multimodal fusion. For example, liu2023multimodal propose MGCL-RA, which presents a contrastive loss of cross-modality to capture similarities between visual and textual modalities. lin2022modeling presente a hierarchical graph contrastive learning (HGraph-CL) framework for multimodal sentiment analysis. It explores sentiment relations across different modalities, employing a graph contrastive learning strategy at the inter-modal level.
@@ 4.1.2 Joint Multimodal Learning
Inter-modal Correlation Modeling. By integrating multimodal graphs into a shared representation space, the joint learning process $Join(\cdot)$ can effectively model inter-modal correlations. saqur2020multimodal proposed a multimodal graph network (MGN) architecture that includes a graph parser and a graph matcher. In the graph parser, they employed a two-layer graph isomorphism network to learn from two multimodal graphs: textual graphs and visual graphs. Subsequently, in the graph matcher, a shared graph neural network is utilized to learn from the multimodal representations. wang2023hypergraph design a unified multimodal phenotypic graph diffusion method to process multimodal brain graphs, capturing inter-modal correlations between different brain graphs (e.g., brain structural graphs and brain functional graphs). Some works focus on dynamic graphs. For example, Melody-GCN [liu2024multimodal] is a multimodal dynamic graph convolutional network designed for stock price prediction tasks. This model specifically addresses three types of graphs, including inter-industry graphs, inter-company graphs, and inter-day graphs. Melody-GCN introduces a multimodal fusing-diffusing block to effectively fuse multimodal representations, enhancing the model’s ability to capture temporal dependencies and interactions.
@@
With a strong relation modeling ability, such as cross-modality relation modeling, the attention mechanism is particularly representative for joint multimodal learning. kim2020hypergraph propose a hypergraph attention network for integrating symbolic information extracted from multimodal symbolic graphs, specifically textual and visual graphs. They developed a semantic similarity measurement-based co-attention mapping function to facilitate the fusion of multimodal information. Graph transformers have gained significant attention because of their ability to effectively capture complex dependencies within graphs. Consequently, many studies have applied multimodal graph transformers to achieve joint multimodal learning. For example, GTGL [cai2022graph] is a graph transformer geometric learning framework to learn on multimodal brain graphs. GTGL implements a graph transformer network with the multi-head self-attention mechanism to jointly learn structural and functional brain graphs. he2023multimodal develop a multimodal graph transformer that incorporates a graph-involved, plug-and-play quasi-attention mechanism for multimodal fusion. Their model operates on three types of graphs, including textual graphs, semantic graphs, and dense region graphs. Specifically, they set a cross-modal mask as an all-zero matrix to learn cross-attention between multimodal features. Moreover, MuSe-GNN [liu2024muse], which features a cross-graph Transformer architecture, is proposed for multimodal biological data integration from gene graphs. In the cross-graph Transformer, a graph neural network is integrated with a multi-head self-attention module, incorporating co-expression information to obtain multimodal gene representations. This design enhances the ability to capture relationships across different gene modalities and improves the quality of multimodal biological data fusion.
@@ 4.2.1 Learning on Multimodal Attributes
$\centering\mathcal{G}=\bigg{(}\mathcal{V},\mathcal{E},\{\mathbf{x}_{i}\}^{|\mathcal{V}|}_{i=1}\bigg{)}.\@add@centering$ (4.3)
@@
Multimodal graph convolutional networks are typical models for learning on multimodal attributes. For example, lian2023gcnet propose GCNet (a graph convolutional network-based graph complete model) for the conversation completion task. In GCNet, multimodal graphs are constructed based on conversation contents, incorporating modalities such as text and images. Each node in the graph contains multimodal features, though some features may be incomplete due to missing modalities. Then, the authors used a relational graph convolutional network to aggregate neighborhood information, enabling more advanced conversation understanding despite incomplete data. 9810283 propose a multicenter and multichannel pooling-based graph convolutional network for multimodal brain graph learning. For each brain region, they integrate information extracted from multimodal brain imaging data and use it as the feature vector for the brain region. A multimodal brain graph is then constructed, where nodes represent different brain regions. To effectively process this graph, the authors introduce a multicenter attention mechanism and multichannel pooling within the graph convolutional network. More recently, meng2024masked propose a method to fuse multimodal features into node attributes. They present an MGLRA (masked graph learning with a recurrent alignment) model, which aims to iteratively refine unimodal representations from multiple modalities. The model first employs a graph convolutional network with a random masking mechanism to facilitate multimodal fusion. Particularly, a cross-modal multi-head attention mechanism is then introduced to learn from the constructed graphs, effectively capturing interactive semantic information across modalities.
@@
Transformer-based and contrastive learning-based methods also attract great attention for learning on multimodal attributes. MarT [000100LDC23] is a Transformer model for multimodal analogical reasoning over knowledge graphs. The authors construct a multimodal knowledge graph (MarKG), where each entity possesses multimodal attributes. This enables the model to reason over relations and analogies between entities across different modalities. chen2022multi design an MSNEA (multimodal siamese network for entity alignment) framework for entity alignment in multimodal knowledge graphs. They first fused entity features from three different modalities, including visual, relational, and attribute modalities, to gain multimodal entity representations. Subsequently, they present a multimodal contrastive learning module to achieve inter-modal enhancement fusion and avoid the overwhelming impact of weak modalities. To address the multimodal emotion recognition task, li2023joyful propose a joint modality fusion-based graph contrastive learning method. Specifically, they design a multimodal fusion mechanism to construct multimodal graphs, where each node contains both unimodal-specific features and global contextual features, enabling a comprehensive information integration from different modalities.
@@
In addition, some other graph learning models are widely proposed for learning on multimodal attributes. TianZGMMC22 introduce a multimodal recipe representation learning model, Recipe2Vec, which leverages a heterogeneous graph neural network to embed multimodal information. They construct heterogeneous recipe graphs by fusing visual, textual, and relational information as node attributes. DGVAE [zhou2024disentangled] converts the visual and textual information of items into consistent embeddings, which are then used to construct item-item graphs where items are represented as nodes. To reconstruct the users’ rating matrix and their textual preferences, DGVAE inputs the word vectors and rating vectors into the graph encoder independently. This approach captures multimodal information to enhance the effectiveness of multimodal recommendations. 3D-HaarFrame [li2024multimodal] is a multimodal graph learning model designed for student engagement prediction. It captures inter-modal correlations within textual, visual, and acoustic data by embedding the multimodal information into a unified graph. The model then applies spectrum-based graph learning to decompose the graph into a set of coefficients at different scales and orientations, efficiently compacting the multimodal representations for improved prediction accuracy.
@@ 4.2.2 Learning on Multimodal Nodes
Knowledge Graphs with Multimodal Nodes. Studies focusing on knowledge graphs with multimodal nodes are rising as a prominent research topic. For example, song2024scene construct a SceneMMKG (scene-driven multimodal knowledge graph), which incorporates both textual and visual nodes. This graph serves as input to a graph convolutional network to facilitate effective learning and representation. wang2020multimodal propose the MKGCN (multimodal knowledge graph convolutional network) model, focusing on knowledge graphs derived from articles containing both text and images. Nodes represent entities extracted from either text or images. MKGCN utilizes two graph convolutional layers to aggregate node features. This approach allows for the generation of high-quality article representations by effectively modeling the connections within text-text, image-image, and text-image relations. Moreover, lyu2021mdnn propose the MDNN (multimodal deep neural network) model, which incorporates a graph neural network module to learn from multimodal drug knowledge graphs for drug-drug interaction prediction. In this knowledge graph, nodes represent drugs and multimodal targets such as molecules and proteins, enabling the model to capture complex interactions between various drug-related modalities. zeng2023multi construct multimodal knowledge hypergraphs through multimodal instance bagging and concept aggregation. The multimodal instance bagging module selects instances by leveraging higher-order relations of hyperedges. A mixed instance bag, containing both images and texts, enhances the diversity of multi-semantic representations. The authors then employ a hypergraph convolutional layer to encode the entities within the graph.
@@
Likewise, many works also focus on other types of graphs with multimodal nodes. For instance, wen2022graph focus on multimodal single-cell graph learning. They introduced scMoGNN, a graph neural network framework that incorporates a cell-feature graph convolution with task-specific heads to analyze multimodal single-cell data across various tasks, such as modality prediction, modality matching, and joint embedding. The model is designed to capture the high-order structural information between cells and modalities. hu2023graph design a graph convolutional network-based graph fusion neural network model to enhance Freezing of Gait (FoG) detection by encoding multimodal information, such as footstep pressure maps and video recordings, into graph nodes. Initially, the features from each modality are encoded as node-level inputs, and a multimodal adjacency matrix is constructed to measure correlations across modalities. Subsequently, They applied graph convolutions to these inputs, focusing on modeling cross-modality relations. SMGCN [kim2024self] is a self-supervised multimodal graph convolutional network designed to operate on multimodal user-item graphs for recommendation tasks. Specifically, the user-item graphs, where edges between multimodal nodes capture cross-modal user preferences, are generated through collaborative multimodal alignment (CMA) and multimodal consistency regularization (MCR). tu2024adaptive proposed AdaIGN (adaptive interactive graph network) for multimodal conversational emotion detection. They emphasized both intra-modal and inter-modal interactions by dynamically selecting nodes and edges across modalities. To achieve this, they employed the Gumbel Softmax trick for adaptive multimodal node selection.
@@
The goal of generative graph learning [guo2022systematic, liu2023generative, cao2024survey, li2025generative] is to understand and learn the distribution of existing graph data, enabling the generation of high-quality graph samples. Generative Graph Learning is significant in studying the underlying structural relationships within graph data, understanding patterns, associations, and hidden information within existing data. Depending on whether auxiliary information is required for graph generation, generative graph learning methods can be categorized into the following two main types: unconditional generative graph learning and conditional generative graph learning.
@@ 5.1.1 One-Shot Generative Graph Learning
Adjacency-matrix-based
@@
(1) MLP-Based Methods These methods use simple MLPs [polsterl2021adversarial, zahirnia2024neural] to construct the graph decoder $g(\textbf{z})$. As shown in Figure 5.1, during the inference phase, a latent graph representation z, which is sampled from a statistical distribution $p(\textbf{z})$, is passed through an MLP-based decoder $g(\textbf{z})$, which generates the adjacency matrix A. Depending on the task requirements, node features F and edge features E may also be generated simultaneously.
@@
(2) Message-Passing-Based These methods generate graphs by iteratively optimizing the initial graph using Message Passing Neural Networks (MPNN) [gilmer2017neural]. As shown in Figure 5.3, an initial adjacency matrix $\textbf{A}_{0}$ and initial node latent representations $\textbf{H}_{0}$ are generated first using the sampled latent representation z.
@@
(3) Flow-Based As shown in Figure 5.3, flow-based methods [madhawa2019graphnvp] generate the graph $\mathcal{G}=(\textbf{A},\textbf{F})$ and the sampled latent prior z through a unique invertible function. Specifically, flow-based methods aim to learn a series of invertible functions $z=f_{L}\circ...\circ f_{1}(G)$ that convert graph data $\mathcal{G}$ into a latent space representation $\textbf{z}\sim P(\textbf{Z})$ with the same dimensionality. During the inference phase, the model uses the inverse transformations $\mathcal{G}=f^{-1}_{1}\circ...\circ f^{-1}_{L}(\textbf{z})$ to reconstruct and generate $\mathcal{G}$ from the latent representation sampled from $\textbf{z}\sim P(\textbf{Z})$.
@@
(4) Diffusion-Based In recent years, diffusion models [cao2024survey] have rapidly emerged as a new generative paradigm and are widely applied across various fields. The working mechanism of diffusion models is divided into two stages: forward diffusion and reverse diffusion. In the forward diffusion stage, the model starts from a simple initial distribution and progressively adds noise, introducing a structured amount of noise at each step. In the reverse diffusion stage, the model begins with noisy data samples and gradually removes the noise, recovering the original data.
@@
Edge-List-Based
@@
(1) Random-Walk-Based In random-walk-based methods, the edge probabilities are estimated by counting the occurrences of edges across numerous random walks, which are sampled from a learned distribution of the graph. NetGAN [bojchevski2018netgan] is introduced to learn the random walk distribution on an observed graph, and then generate a set of random walks.
@@
Inspired by NetGAN, subsequent research further refines random walks to better suit various tasks. For example, the temporal random walk is used in generative temporal graph learning [gupta2022tigger]. Similarly, TIARA [lee2023time] utilize the time-aware random walks to enhance the dynamic graph.
@@
In addition to the two categories of research mentioned above, there are also other edge-list-based approaches for graph generation. For instance, a few studies [jang2023simple] propose edge-list-based representation methods to achieve lightweight graph generation. These approaches often use Depth-First Search (DFS) or Breadth-First Search (BFS) to construct edges, ensuring unique node ordering while maintaining high scalability. Additionally, luo2023fast focus on the global information implied by the graph’s spectrum. By learning and generating the graph’s eigenvalues and eigenvectors, these methods aim to capture the shape information of the graph, leading to an improved generation performance.

### v2

@@ 2.3.2 Mini-batch Training
Despite their effectiveness, typical graph learning methods face challenges when applied to dynamic or evolving graphs. Static models assume that the graph structure and node features do not change over time, which is often unrealistic in real-world scenarios where graph structures continuously evolve [barros2021survey_dynamicgraphsurvey, gravina2024deep_dynamicgraphSurveyBenchmark, febrinanto2023graph_GLL, zheng2025survey]. Such phenomena are common across many domains, prompting researchers to propose a variety of datasets and benchmarks for further investigation.
@@ 3.2 Spatiotemporal Graph Learning
$A_{u,v}=\frac{\exp\left(\frac{\log\widetilde{A}_{u,v}+\epsilon_{v}}{\tau}\right)}{\sum_{w\in\mathcal{N}_{u}}\exp\left(\frac{\log\widetilde{A}_{u,v}+\epsilon_{w}}{\tau}\right)}.$ (3.2)
@@ 3.2.3 Attention-based Methods
Data in real-world scenarios often come from different sources. Multimodal fusion aims to integrate information from diverse data sources, such as visual, textual, and acoustic data. However, multimodal data often exhibits intricate relational dependencies, including both inter-modal and intra-modal correlations. Traditional multimodal learning methods that rely on basic modality alignment and simple fusion techniques often fail to effectively capture and integrate the rich, complementary information across different modalities. To address this challenge, multimodal graph learning has emerged as a powerful approach [ektefaie2023multimodal, peng2024learning, xie2024multimodal, he2025unigraph2]. It leverages graph structures to model and learn the intricate inter-modal and intra-modal correlations, enabling a more nuanced integration of multimodal data.
@@ 4.1.1 Separate Unimodal Learning
Modality-specific Representation. Separate unimodal learning exhibits the advantage of capturing modality-specific representation. Specifically, the modality-specific representation $\mathbf{Z}_{i}$ for the modality $i$ can be effectively extracted using various independent graph learning models. For example, MGAT [tao2020mgat] applies embedding propagation layers to perform message passing on each single-modal user item graph, capturing user preferences on each modality. guo2024lgmrec proposed local and global graph learning-guided multimodal recommender (LGMRec) for learning on multimodal user-item graphs. It introduces a local graph embedding module to perform message propagation on textual and visual user-item graphs separately, capturing intra-modal local user interests and generating modality-specific representations. MKGformer [chen2022hybrid] is a hybrid Transformer model designed for the multimodal knowledge graph completion task. In MKGformer, two types of knowledge graph are constructed: visual knowledge graphs and textual knowledge graphs. A visual Transformer and a textual Transformer, both equipped with self-attention modules, are then applied to these knowledge graphs to learn modality-specific representations.
@@
  • Contrastive learning-based fusion: Some works present multimodal graph contrastive learning frameworks. These frameworks apply a contrastive objective strategy to capture inter-modal and intra-modal similarities and differences, enabling multimodal fusion. For example, liu2023multimodal proposed MGCL-RA, which presents a contrastive loss of cross-modality to capture similarities between visual and textual modalities. lin2022modeling presented a hierarchical graph contrastive learning (HGraph-CL) framework for multimodal sentiment analysis. It explores sentiment relations across different modalities, employing a graph contrastive learning strategy at the inter-modal level.
@@ 4.1.2 Joint Multimodal Learning
Inter-modal Correlation Modeling. By integrating multimodal graphs into a shared representation space, the joint learning process $Join(\cdot)$ can effectively model inter-modal correlations. saqur2020multimodal proposed a multimodal graph network (MGN) architecture that includes a graph parser and a graph matcher. In the graph parser, they employed a two-layer graph isomorphism network to learn from two multimodal graphs: textual graphs and visual graphs. Subsequently, in the graph matcher, a shared graph neural network is utilized to learn from the multimodal representations. wang2023hypergraph designed a unified multimodal phenotypic graph diffusion method to process multimodal brain graphs, capturing inter-modal correlations between different brain graphs (e.g., brain structural graphs and brain functional graphs).
Some works focus on dynamic graphs. For example, Melody-GCN [liu2024multimodal] is a multimodal dynamic graph convolutional network designed for stock price prediction tasks. This model specifically addresses three types of graphs, including inter-industry graphs, inter-company graphs, and inter-day graphs. Melody-GCN introduces a multimodal fusing-diffusing block to effectively fuse multimodal representations, enhancing the model’s ability to capture temporal dependencies and interactions.
@@
With a strong relation modeling ability, such as cross-modality relation modeling, the attention mechanism is particularly representative for joint multimodal learning. kim2020hypergraph proposed a hypergraph attention network for integrating symbolic information extracted from multimodal symbolic graphs, specifically textual and visual graphs. They developed a semantic similarity measurement-based co-attention mapping function to facilitate the fusion of multimodal information. Graph transformers have gained significant attention because of their ability to effectively capture complex dependencies within graphs. Consequently, many studies have applied multimodal graph transformers to achieve joint multimodal learning. For example, GTGL [cai2022graph] is a graph transformer geometric learning framework to learn on multimodal brain graphs. GTGL implements a graph transformer network with the multi-head self-attention mechanism to jointly learn structural and functional brain graphs. he2023multimodal developed a multimodal graph transformer that incorporates a graph-involved, plug-and-play quasi-attention mechanism for multimodal fusion. Their model operates on three types of graphs, including textual graphs, semantic graphs, and dense region graphs. Specifically, they set a cross-modal mask as an all-zero matrix to learn cross-attention between multimodal features. Moreover, MuSe-GNN [liu2024muse], which features a cross-graph Transformer architecture, is proposed for multimodal biological data integration from gene graphs. In the cross-graph Transformer, a graph neural network is integrated with a multi-head self-attention module, incorporating co-expression information to obtain multimodal gene representations. This design enhances the ability to capture relationships across different gene modalities and improves the quality of multimodal biological data fusion.
@@ 4.2.1 Learning on Multimodal Attributes
[…]
@@
Multimodal graph convolutional networks are typical models for learning on multimodal attributes. For example, lian2023gcnet proposed GCNet (a graph convolutional network-based graph complete model) for the conversation completion task. In GCNet, multimodal graphs are constructed based on conversation contents, incorporating modalities such as text and images. Each node in the graph contains multimodal features, though some features may be incomplete due to missing modalities. Then, the authors used a relational graph convolutional network to aggregate neighborhood information, enabling more advanced conversation understanding despite incomplete data. 9810283 propose a multicenter and multichannel pooling-based graph convolutional network for multimodal brain graph learning. For each brain region, they integrated information extracted from multimodal brain imaging data and used it as the feature vector for the brain region. A multimodal brain graph is then constructed, where nodes represent different brain regions. To effectively process this graph, the authors introduce a multicenter attention mechanism and multichannel pooling within the graph convolutional network. More recently, meng2024masked proposed a method to fuse multimodal features into node attributes. They present an MGLRA (masked graph learning with a recurrent alignment) model, which aims to iteratively refine unimodal representations from multiple modalities. The model first employs a graph convolutional network with a random masking mechanism to facilitate multimodal fusion. Particularly, a cross-modal multi-head attention mechanism is then introduced to learn from the constructed graphs, effectively capturing interactive semantic information across modalities.
@@
Transformer-based and contrastive learning-based methods also attract great attention for learning on multimodal attributes. MarT [000100LDC23] is a Transformer model for multimodal analogical reasoning over knowledge graphs. The authors constructed a multimodal knowledge graph (MarKG), where each entity possesses multimodal attributes. This enables the model to reason over relations and analogies between entities across different modalities. chen2022multi designed an MSNEA (multimodal siamese network for entity alignment) framework for entity alignment in multimodal knowledge graphs. They first fused entity features from three different modalities, including visual, relational, and attribute modalities, to gain multimodal entity representations. Subsequently, they present a multimodal contrastive learning module to achieve inter-modal enhancement fusion and avoid the overwhelming impact of weak modalities. To address the multimodal emotion recognition task, li2023joyful proposed a joint modality fusion-based graph contrastive learning method. Specifically, they designed a multimodal fusion mechanism to construct multimodal graphs, where each node contains both unimodal-specific features and global contextual features, enabling a comprehensive information integration from different modalities.
@@
In addition, some other graph learning models are widely proposed for learning on multimodal attributes. TianZGMMC22 introduced a multimodal recipe representation learning model, Recipe2Vec, which leverages a heterogeneous graph neural network to embed multimodal information. They constructed heterogeneous recipe graphs by fusing visual, textual, and relational information as node attributes. DGVAE [zhou2024disentangled] converts the visual and textual information of items into consistent embeddings, which are then used to construct item-item graphs where items are represented as nodes. To reconstruct the users’ rating matrix and their textual preferences, DGVAE inputs the word vectors and rating vectors into the graph encoder independently. This approach captures multimodal information to enhance the effectiveness of multimodal recommendations. 3D-HaarFrame [li2024multimodal] is a multimodal graph learning model designed for student engagement prediction. It captures inter-modal correlations within textual, visual, and acoustic data by embedding the multimodal information into a unified graph. The model then applies spectrum-based graph learning to decompose the graph into a set of coefficients at different scales and orientations, efficiently compacting the multimodal representations for improved prediction accuracy.
@@ 4.2.2 Learning on Multimodal Nodes
Knowledge Graphs with Multimodal Nodes. Studies focusing on knowledge graphs with multimodal nodes are rising as a prominent research topic. For example, song2024scene constructed a SceneMMKG (scene-driven multimodal knowledge graph), which incorporates both textual and visual nodes. This graph serves as input to a graph convolutional network to facilitate effective learning and representation. wang2020multimodal proposed the MKGCN (multimodal knowledge graph convolutional network) model, focusing on knowledge graphs derived from articles containing both text and images. Nodes represent entities extracted from either text or images. MKGCN utilizes two graph convolutional layers to aggregate node features. This approach allows for the generation of high-quality article representations by effectively modeling the connections within text-text, image-image, and text-image relations. Moreover, lyu2021mdnn proposed the MDNN (multimodal deep neural network) model, which incorporates a graph neural network module to learn from multimodal drug knowledge graphs for drug-drug interaction prediction. In this knowledge graph, nodes represent drugs and multimodal targets such as molecules and proteins, enabling the model to capture complex interactions between various drug-related modalities. zeng2023multi constructed multimodal knowledge hypergraphs through multimodal instance bagging and concept aggregation. The multimodal instance bagging module selects instances by leveraging higher-order relations of hyperedges. A mixed instance bag, containing both images and texts, enhances the diversity of multi-semantic representations. The authors then employ a hypergraph convolutional layer to encode the entities within the graph.
@@
Likewise, many works also focus on other types of graphs with multimodal nodes. For instance, wen2022graph focused on multimodal single-cell graph learning. They introduced scMoGNN, a graph neural network framework that incorporates a cell-feature graph convolution with task-specific heads to analyze multimodal single-cell data across various tasks, such as modality prediction, modality matching, and joint embedding. The model is designed to capture the high-order structural information between cells and modalities. hu2023graph designed a graph convolutional network-based graph fusion neural network model to enhance Freezing of Gait (FoG) detection by encoding multimodal information, such as footstep pressure maps and video recordings, into graph nodes. Initially, the features from each modality are encoded as node-level inputs, and a multimodal adjacency matrix is constructed to measure correlations across modalities. Subsequently, They applied graph convolutions to these inputs, focusing on modeling cross-modality relations. SMGCN [kim2024self] is a self-supervised multimodal graph convolutional network designed to operate on multimodal user-item graphs for recommendation tasks. Specifically, the user-item graphs, where edges between multimodal nodes capture cross-modal user preferences, are generated through collaborative multimodal alignment (CMA) and multimodal consistency regularization (MCR). tu2024adaptive proposed AdaIGN (adaptive interactive graph network) for multimodal conversational emotion detection. They emphasized both intra-modal and inter-modal interactions by dynamically selecting nodes and edges across modalities. To achieve this, they employed the Gumbel Softmax trick for adaptive multimodal node selection.
@@
The goal of generative graph learning [guo2022systematic, liu2023generative, cao2024survey, li2025generative] is to understand and learn the distribution of existing graph data, enabling the generation of high-quality graph samples. Generative Graph Learning is significant in studying the underlying structural relationships within graph data, understanding patterns, associations, and hidden information within existing data. To this end, researchers have proposed a large number of synthetic datasets and real-world datasets for research purposes.
Depending on whether auxiliary information is required for graph generation, generative graph learning methods can be categorized into the following two main types: unconditional generative graph learning and conditional generative graph learning.
@@ 5.1.1 One-Shot Generative Graph Learning
Adjacency-matrix-based Methods
@@
(1) MLP-Based Methods. These methods use simple MLPs [polsterl2021adversarial, zahirnia2024neural] to construct the graph decoder $g(\textbf{z})$. As shown in Figure 5.1, during the inference phase, a latent graph representation z, which is sampled from a statistical distribution $p(\textbf{z})$, is passed through an MLP-based decoder $g(\textbf{z})$, which generates the adjacency matrix A. Depending on the task requirements, node features F and edge features E may also be generated simultaneously.
@@
(2) Message-Passing-Based Methods. These methods generate graphs by iteratively optimizing the initial graph using Message Passing Neural Networks (MPNN) [gilmer2017neural]. As shown in Figure 5.2, an initial adjacency matrix $\textbf{A}_{0}$ and initial node latent representations $\textbf{H}_{0}$ are generated first using the sampled latent representation z.
@@
(3) Flow-Based Methods. As shown in Figure 5.3, flow-based methods [madhawa2019graphnvp] generate the graph $\mathcal{G}=(\textbf{A},\textbf{F})$ and the sampled latent prior z through a unique invertible function. Specifically, flow-based methods aim to learn a series of invertible functions $z=f_{L}\circ...\circ f_{1}(G)$ that convert graph data $\mathcal{G}$ into a latent space representation $\textbf{z}\sim P(\textbf{Z})$ with the same dimensionality. During the inference phase, the model uses the inverse transformations $\mathcal{G}=f^{-1}_{1}\circ...\circ f^{-1}_{L}(\textbf{z})$ to reconstruct and generate $\mathcal{G}$ from the latent representation sampled from $\textbf{z}\sim P(\textbf{Z})$.
@@
(4) Diffusion-Based Methods. In recent years, diffusion models [cao2024survey] have rapidly emerged as a new generative paradigm and are widely applied across various fields. The working mechanism of diffusion models is divided into two stages: forward diffusion and reverse diffusion. In the forward diffusion stage, the model starts from a simple initial distribution and progressively adds noise, introducing a structured amount of noise at each step. In the reverse diffusion stage, the model begins with noisy data samples and gradually removes the noise, recovering the original data.
@@
Edge-List-Based Methods.
@@
(1) Random-Walk-Based Methods. In random-walk-based methods, the edge probabilities are estimated by counting the occurrences of edges across numerous random walks, which are sampled from a learned distribution of the graph. NetGAN [bojchevski2018netgan] is introduced to learn the random walk distribution on an observed graph, and then generate a set of random walks.
@@
Inspired by NetGAN, subsequent research further refines random walks to better suit various tasks. For example, the temporal random walk is used in generative temporal graph learning [gupta2022tigger]. Similarly, TIARA [lee2023time] utilizes the time-aware random walks to enhance the dynamic graph.
@@
In addition to the two categories of research mentioned above, there are also other edge-list-based approaches for graph generation. For instance, a few studies [jang2023simple] propose edge-list-based representation methods to achieve lightweight graph generation. These approaches often use Depth-First Search (DFS) or Breadth-First Search (BFS) to construct edges, ensuring unique node ordering while maintaining high scalability. Additionally, luo2023fast focused on the global information implied by the graph’s spectrum. By learning and generating the graph’s eigenvalues and eigenvectors, these methods aim to capture the shape information of the graph, leading to an improved generation performance.